# MetaGFN: Exploring Distant Modes with Adapted Metadynamics for Continuous GFlowNets

### v1

@@ 5 Limitations
For metadynamics to be an effective sampler, the CVs must be low-dimensional and bounded, properties that were satisfied in both our experimental environments. Therefore, it is necessary to either know such CVs in advance, assuming they exist or learn them automatically from data Sidky et al. (2020). An alternative approach would be to learn CVs adaptively by parameterising the CV function by a neural network and updating its parameters by back-propagating through the GFlowNet loss when training on MetaGFN trajectories. A final improvement could be to replace the metadynamics algorithm itself with a variant with smoother convergence properties, such as well-tempered metadynamics Barducci et al. (2008) or on-the-fly probability enhanced sampling (OPES) Invernizzi (2021). We leave these extensions for future work.
@@ 6 Conclusions
In molecular dynamics, atomic forces can be computed as the gradient of the potential, whereas continuous GFlowNets tackle problems where the reward function is a black box and gradients are inaccessible. We demonstrated how the method could be adapted by updating a kernel density estimate of the reward function on-the-fly, and proved that this is guaranteed to explore the space in an appropriate limit. Our empirical investigations show that MetaGFN offers a computationally efficient means to explore new modes in environments where prior knowledge of collective variables exists. Importantly, this work advocates an approach wherein techniques derived from molecular modelling can be adapted for machine learning tasks. Looking ahead, we anticipate that this could be a fruitful area of cross-disciplinary research, where existing ideas from the enhanced sampling literature can find applications in a broad range of generative modelling and reinforcement learning tasks.
@@ Appendix B Langevin dynamics
In the above, $x,p\in\mathbb{R}^{D}$ are vectors of instantaneous position and momenta respectively, $F\vcentcolon\mathbb{R}^{D}\rightarrow\mathbb{R}^{D}$ is a force function, $W(t)$ is a vector of $D$ independent Wiener processes, $M\in\mathbb{R}^{D}\times\mathbb{R}^{D}$ is a constant diagonal mass matrix, and $\gamma,\beta>0$ are constant scalars which can be interpreted as a friction coefficient and inverse temperature respectively. In conventional Langevin dynamics, the force function is given by the gradient of the potential energy function, $F=-\nabla V$, where $V\vcentcolon\mathbb{R}^{D}\rightarrow\mathbb{R}$ and the dynamics are ergodic with respect to the Gibbs-Boltzmann density
@@
where $H(x,p)=p^{T}M^{-1}p/2+V(x)$ is the Hamiltonian. Since the Hamiltonian is separable in position and momenta terms, the marginal Gibbs-Boltzmann density is position space is simply $\rho_{\beta}(x)\propto e^{-\beta V(x)}$.
@@ Appendix C Proofs
First we take the $t\rightarrow\infty$ limit. Since the log function is continuous, the limit and log can be interchanged and we have
$\lim_{t\rightarrow\infty}\hat{V}(x,t)=-\frac{1}{\beta^{\prime}}\log\left(\lim_{t\rightarrow\infty}\left(\frac{\hat{R}(x,t)}{\hat{N}(x,t)+\epsilon}\right)+\epsilon\right).$
Recall the uniform time discretisation $t_{n}=n\Delta t$ of metadynamics (Algorithm 1). Thus, we can write
$\hat{R}(x,t_{n})=\sum_{i=1}^{n}r(x_{i})\exp{\left(-\sum_{i=1}^{d}\frac{(z_{i}(x)-z(x_{i}))^{2}}{2{\sigma}_{i}^{2}}\right)},$ (8)
$\hat{N}(x,t_{n})=\sum_{i=1}^{n}\exp{\left(-\sum_{i=1}^{d}\frac{(z_{i}(x)-z(x_{i}))^{2}}{2{\sigma}_{i}^{2}}\right)}.$ (9)
Since the domain is bounded, we know that for fixed $x$, both (8) and (9) have limit at infinity, i.e. $\lim_{t\rightarrow\infty}\hat{R}(x,t)=\infty$ and $\lim_{t\rightarrow\infty}\hat{N}(x,t)=\infty$. Hence, by Lemma C.1, we have
$\lim_{t\rightarrow\infty}\frac{\hat{R}(x,t)}{\hat{N}(x,t)+\epsilon}=\lim_{t\rightarrow\infty}\frac{\hat{R}(x,t)}{\hat{N}(x,t)},$
@@
$\lim_{t\rightarrow\infty}\frac{\hat{R}(x,t)}{\hat{N}(x,t)}=\lim_{n\rightarrow\infty}\frac{\hat{R}(x,t_{n})}{\hat{N}(x,t_{n})}=\lim_{n\rightarrow\infty}\frac{\sum_{i=1}^{n}r(x_{i})\exp\left(-\sum_{i=1}^{d}\frac{(z_{i}-z(x_{i}))^{2}}{2{\sigma}_{i}^{2}}\right)}{\sum_{i=1}^{n}\exp\left(-\sum_{i=1}^{d}\frac{(z_{i}-z(x_{i}))^{2}}{2{\sigma}_{i}^{2}}\right)}.$
Recall that metadynamics eventually leads to uniform sampling over the domain, independent of the potential. Hence, since $R$ and $z(x)$ are analytic, by Lemma C.2 we may replace the metadynamics samples $x_{i}$ with samples from a uniform distribution, denoted $u_{i}$:
$\lim_{n\rightarrow\infty}\frac{\hat{R}(x,t_{n})}{\hat{N}(x,t_{n})}=\lim_{n\rightarrow\infty}\frac{\sum_{i=1}^{n}r(u_{i})\exp\left(-\sum_{i=1}^{d}\frac{(z_{i}-z(u_{i}))^{2}}{2{\sigma}_{i}^{2}}\right)}{\sum_{i=1}^{n}\exp\left(-\sum_{i=1}^{d}\frac{(z_{i}-z(u_{i}))^{2}}{2{\sigma}_{i}^{2}}\right)}.$
@@
$\lim_{n\rightarrow\infty}\frac{\hat{R}(x,t_{n})}{\hat{N}(x,t_{n})}=\frac{\int_{D}r(x^{\prime})\exp\left(-\sum_{i=1}^{d}\frac{(z_{i}-z(x^{\prime}_{i}))^{2}}{2{\sigma}_{i}^{2}}\right)\mathrm{d}x^{\prime}}{\int_{D}\exp\left(-\sum_{i=1}^{d}\frac{(z_{i}-z(x^{\prime}_{i}))^{2}}{2{\sigma}_{i}^{2}}\right)\mathrm{d}x^{\prime}},$
where $D\subset\mathbb{R}^{d}$ is the domain. The limit is therefore a (scaled) convolution of the reward function with a Gaussian in the collective variable space with width vector $\sigma$. Taking the limit $\sigma_{i}\rightarrow 0$ for all $i\in\left\{1,2,\dots,d\right\}$, the Gaussian convergences to a delta distribution in the collective variable space and we have
$\lim_{\sigma\rightarrow 0}\lim_{n\rightarrow\infty}\frac{\hat{R}(x,t_{n})}{\hat{N}(x,t_{n})}=\int_{D}r(x^{\prime})\delta(z-z(x^{\prime}))\mathrm{d}x^{\prime}.$
@@
$\lim_{\epsilon\rightarrow 0}\lim_{\sigma\rightarrow 0}\lim_{t\rightarrow\infty}\hat{V}(x,t)$ $=-\frac{1}{\beta^{\prime}}\lim_{\epsilon\rightarrow 0}\log\left(\int_{D}r(x^{\prime})\delta(z-z(x^{\prime}))\mathrm{d}x^{\prime}+\epsilon\right)$ (10)
$=-\frac{1}{\beta^{\prime}}\int_{D}\log\left(r(x^{\prime})\right)\delta(z-z(x^{\prime}))\mathrm{d}x^{\prime}$ (11)
$=\int_{D}V(x^{\prime})\delta(z-z(x^{\prime}))\mathrm{d}x^{\prime}\vcentcolon=V(z),$ (12)
@@
Appendix D Algorithms
Input : Current state $(x_{t},p_{t})$. Force $F$.
Output : State $(x_{t+\Delta t},p_{t+\Delta t})$ at the next timestep.
1
3 $x_{t+\Delta t}=x_{t}+p_{t}\Delta t$
5return $(x_{t+\Delta t},p_{t+\Delta t})$
Algorithm 2 Euler-Maruyama Langevin Dynamics Step
In the algorithm below, we present the variant of MetaGFN where we store Adaptive Metadynamics samples in the replay buffer and regenerate trajectories using the current backward policy when retrieving from the replay buffer. This is the variant we used in our experiments in the main text.
Input : Forward policy $P_{F}$. Backwards policy $P_{B}$. Loss function $L$.
Parameters : How often to run Adaptive Metadynamics batches, freqMD. How often to run replay buffer batches, freqRB. Batch size, $b$. Stride, $n\in\mathbb{Z}^{+}$. Time step, $\Delta t>0$
1
2for each episode do:
3if episode number is divisible by freqMD:
4   Run Adaptive Metadynamics (batch size $b$) for time $n\Delta t$, obtain samples $\{x_{1},\dots,x_{b}\}$
5   Push $\{x_{1},\dots,x_{b}\}$ to the replay buffer
7elif episode number is divisible by freqRB:
8   Random sample $\{x_{1},\dots,x_{b}\}$ from the replay buffer
10
11else:
12   Generate trajectories $\{\tau_{1},\dots,\tau_{b}\}$ on-policy
13
14 Compute loss $l=\sum_{i=1}^{b}L(\tau_{i},P_{F},P_{B})$
 Take gradient step on loss $l$
Algorithm 3 MetaGFN
Appendix E Experiment details: line environment
E.1 Experimental setup
Parameterisation We parameterise $\hat{p}_{F}$, $\hat{p}_{B}$ and the flow $\hat{f}$ through an MLP with 3 hidden layers, 256 hidden units per layer. We use the GELU activation function and dropout probability $0.2$ after each layer. This defines the torso of the MLP. Connecting from this common torso, the MLP has three single-layer, fully connected heads. The first two heads have output dimension $9$ and parameterise the $3$ means ($\mu$), standard deviations ($\sigma$) and weights ($w$) of the mixture of Gaussians for the forward and backward policies respectively. The third head has output dimension $1$ and parameterises the flow function $\hat{f}$. The mean and standard deviation outputs are passed through a sigmoid function and transformed so that they map to the ranges $\mu\in(-14,14)$ and $\sigma\in(0.1,1)$. The mixture weights are normalised with the softmax function. The exception to this parameterisation is the backward transition to the source state which is fixed to be the Dirac delta distribution centred on the source, i.e. $\hat{p}_{B}(s_{0}|s_{1};\theta)=\delta_{s_{0}}$. For the TB loss, we treat $\log Z_{\theta}$ as a separate learnable parameter.
Replay buffer The replay buffer has capacity for $10^{4}$ trajectories. Trajectories are stored in the replay buffer only if the terminal state’s reward exceeds $10^{-3}$. When drawing a replay buffer batch, trajectories are bias-sampled: $50\%$ randomly drawn from the upper $30\%$ of trajectories with the highest rewards, and the remaining $50\%$ randomly drawn from the lower $70\%$.
@@ E.1 Experimental setup
Noisy exploration Noisy exploration is defined by adding an additional constant, $\bar{\sigma}$, to the standard deviations of the Gaussian distributions of the forward and backward policies. Specifically, the forward policy becomes $\hat{p}_{F}(s_{t},s_{t-1};\theta)=\sum_{i=1}^{3}w_{i}\mathcal{N}(\mu_{i},(\sigma_{i}+\bar{\sigma})^{2})$, and similarly for the backward policy. We schedule the value of $\bar{\sigma}$ so that it decreases during training according to an exponential-flat schedule:
@@
where $j\in(1,\dots,B)$ is the batch number and $\bar{\sigma}_{0}=2$ is the initial noise, plotted in Figure 8.
Thompson Sampling We use $10$ heads with the bootstrapping probability parameter set to $0.3$.
@@
MetaGFN We use $\Delta t=0.05$, $n=2$, $\beta=1$, $\gamma=2$, $w=0.15$, $\sigma=0.1$, $\epsilon=10^{-3}$. The domain of Adaptive Metadynamics is restricted to $[-5,23]$ and reflection conditions are imposed at the boundary. The bias and KDE potentials are stored on a uniform grid with grid spacing $0.01$. Initial metadynamics samples are drawn from a Gaussian distribution, mean 00 and variance $1$, and initial momenta from a Gaussian distribution, mean 00 and variance $0.5$.
Training parameters In all experiments, we use batch size $b=64$ and for $B=10^{5}$ batches. We use a learning rate with a linear schedule, starting at $10^{-3}$ and finishing at 00. For the TB loss, we train the $\log Z_{\theta}$ with a higher initial learning rate of $10^{-1}$ (also linearly scheduled). For the STB loss, we use $\lambda=0.9$, a value that has worked well in the discrete setting Madan et al. [2022]. We use the Adam optimiser with gradient clipping. For all loss functions, we clip the minimum log-reward signal at $-10$. This enables the model to learn despite regions of near-zero reward between the modes of $r(x)$.
@@
Evaluation The L1 error between the known reward distribution, $\rho(x)=r(x)/Z$, and the empirical on-policy distribution, denoted $\hat{\rho}(x)$, estimated by sampling $10^{4}$ on-policy trajectories and computing the empirical distribution over terminal states. Specifically, we compute
$\text{error}=\frac{1}{2}\int_{-5}^{23}\left|\hat{\rho}(x)-\frac{r(x)}{Z}\right|\mathrm{d}x,$ (14)
where the integral is estimated by a discrete sum with grid spacing $0.01$. Note that this error is normalised such that for all valid probability distributions $\hat{\rho}(x)$, we have $0\leq\text{error}\leq 1$.
Compute resources Experiments are performed in PyTorch on a desktop with 32Gb of RAM and a 12-core 2.60GHz i7-10750H CPU. It takes approximately $1$ hour to train a continuous GFlowNet in this environment with $B=10^{5}$ batches. The additional computational expense of running Adaptive Metadynamics was negligible compared to the training time of the models.
@@
E.2 Results
On-policy distributions Figure 9 shows the forward policy and replay buffer distributions (with bias sampling) after training for $10^{5}$ iterations with TB loss. MetaGFN is the only method that can uniformly populate the replay buffer and consistently learn all three peaks.
Adaptive Metadynamics Figure 10(a) shows the L1 error between the density implied by the KDE potential and the true reward distribution during a typical training run. Figure 10(b) shows the resulting $\hat{V}$ and $V_{\text{bias}}$ at the end of the training. By (1), Adaptive Metadynamics has fully explored the central peaks. At (2), the third peak is discovered, prompting rapid adjustment of the KDE potential. By $2.5\times 10^{4}$ iterations, a steady state is reached and the algorithm is sampling the domain uniformly.
@@ E.2 Results
Comparing MetaGFN Variants We consider three MetaGFN variants. The first variant, always backwards sample, regenerates the entire trajectory using the current backward policy when pulling from the replay buffer. The second variant, reuse initial backwards sample, generates the trajectory when first added to the replay buffer and reuses the entire trajectory if subsequently sampled. The third variant, with noise, is to always backwards sample with noisy exploration as per equation (13). We plot the L1 policy errors in Figure 11. We observe that always backward sample is better than reuse initial backwards sample for all loss functions. For DB and TB losses, there is no evidence of any benefit of adding noise, whereas noise improves training for STB loss, performing very similarly to TB loss without noise.
@@
Appendix F Experiment details: grid environment
Parameterisation We parameterise $\hat{p}_{F}$, $\hat{p}_{B}$ and the flow $\hat{f}$ through an MLP with 3 hidden layers, 512 hidden units per layer, GELU activation function and dropout probability $0.2$, similar to the Line Environment. The first two heads have output dimension $24$ and parameterise the $8$ means, $12$ standard deviations and $4$ weights of the mixture of bivariate Gaussians. The third head has output dimension $1$ and parameterises the flow function $\hat{f}$. The means and standard deviations are mapped to the range $\mu\in(-15,15)$ and $\sigma\in(0.1,7)$ through sigmoid functions. Mixture weights are normalised with the softmax function.
Replay buffer Identical to the Line Environment replay buffer, except trajectories are only stored when the terminal state’s reward exceeds $10^{-4}$.
Noisy exploration We use the same noise profile as the Line Environment, equation (13). The noise $\bar{\sigma}$ is added equally to both dimensions of the standard deviation.
Thompson sampling We use $10$ heads with the bootstrap probability parameter set to $0.3$.
MetaGFN We use $\texttt{freqRB}=2$, $\texttt{freqMD}=10$, $\Delta t=0.35$, $n=3$, $\beta=1$, $\gamma=2$, $w=0.10$, $\sigma=2$. The bias and KDE potentials are stored on a uniform grid with grid spacing $0.075$. Initial samples are drawn from a Gaussian distribution, mean centred $(0,0)$, variance $\sigma^{2}=(1,1)$. Momenta are initialised to zero.
Training parameters The same as for the Line Environment, see Appendix E.1.
Evaluation The L1 error between the histogram of the terminal states of $10^{4}$ on-policy samples and the exact distribution is computed. $\text{error}=\frac{1}{2}\int_{-15}^{15}\int_{-15}^{15}\left|\hat{\rho}(x,y)-\frac{r(x,y)}{Z}\right|\mathrm{d}x\mathrm{d}y$, estimated by a discrete sum with grid spacing $0.075$.
Compute resources Experiments are performed in PyTorch on a desktop with 32Gb of RAM and a 12-core 2.60GHz i7-10750H CPU. It takes approximately $10$ hours to train a continuous GFlowNet in this environment with $B=10^{5}$ batches. The additional computational expense of running Adaptive Metadynamics is less than $10\%$.
Appendix G Experiment details: alanine dipeptide environment

### v2

@@ 5 Limitations
For metadynamics to be an effective sampler, the CVs must be low-dimensional and bounded. Either these CVs should be known analytically (as in our experiments) or learnt from data [42]. Alternatively, CVs could be learnt adaptively by parameterising $z(x;\theta)$ by a neural network and updating its parameters by back-propagating through the GFlowNet loss during training. The metadynamics algorithm could also be replaced with a variant with smoother convergence properties, such as well-tempered metadynamics [45] or on-the-fly probability enhanced sampling (OPES) [46]. We leave these as extensions for future work.
@@ 6 Conclusions
In molecular dynamics, atomic forces can be computed as the gradient of the potential, whereas continuous GFlowNets tackle problems where the reward function is a black box and gradients are inaccessible. We demonstrated how the method could be adapted by updating a kernel density estimate of the reward function on-the-fly, and proved that this is guaranteed to explore the space in an appropriate limit. Our empirical investigations show that MetaGFN offers a computationally efficient means to explore new modes in environments where prior knowledge of collective variables exists. Importantly, this work advocates an approach wherein techniques derived from molecular modelling can be adapted for machine learning tasks. Looking ahead, we anticipate that this could be a fruitful area of cross-disciplinary research, where existing ideas from the enhanced sampling literature can find further applications in generative modelling and reinforcement learning.
@@ Appendix B Langevin dynamics
In the above, $x,p\in\mathbb{R}^{D}$ are vectors of instantaneous position and momenta respectively, $F\vcentcolon\mathbb{R}^{D}\rightarrow\mathbb{R}^{D}$ is a force function, $W(t)$ is a vector of $D$ independent Wiener processes, $M\in\mathbb{R}^{D}\times\mathbb{R}^{D}$ is a constant diagonal mass matrix, and $\gamma,\beta>0$ are constant scalars which can be interpreted as a friction coefficient and inverse temperature respectively. In conventional Langevin dynamics, the force function is given by the gradient of the potential energy function, $F(x)=-\nabla V(x)$, where $V\vcentcolon\mathbb{R}^{D}\rightarrow\mathbb{R}$ and the dynamics are ergodic with respect to the Gibbs-Boltzmann density
@@
where $H(x,p)=p^{T}M^{-1}p/2+V(x)$ is the Hamiltonian. Since the Hamiltonian is separable in position and momenta terms, the marginal Gibbs-Boltzmann density is position space is simply $\rho_{\beta}(x)\propto e^{-\beta V(x)}$, often refered to as the Gibbs density.
We present the Euler-Maruyama numerican scheme for integrating equations 3 below. This is called in line 1 of Adapted Metadynamics (Algorithm 1).
Input : Current state $(x_{t},p_{t})$. Force $F$.
Output : State $(x_{t+\Delta t},p_{t+\Delta t})$ at the next timestep.
1
3 $x_{t+\Delta t}=x_{t}+p_{t}\Delta t$
5return $(x_{t+\Delta t},p_{t+\Delta t})$
Algorithm 3 Euler-Maruyama Langevin Dynamics Step
@@ Appendix C Proofs
First we take the $n\rightarrow\infty$ limit. Recall the notation from Section 3, i.e. assume collective variables $z(x)=(z_{1}(x),\dots,z_{d}(x))$ where $z\vcentcolon\mathcal{X}\rightarrow\mathcal{Z}$, and $\mathcal{Z}$ is $d$-dimensional. Since the log function is continuous, the limit and log can be interchanged and we have
$\lim_{n\rightarrow\infty}\hat{V}(z,t_{n})=-\frac{1}{\beta^{\prime}}\log\left(\lim_{n\rightarrow\infty}\left(\frac{\hat{R}(z,t_{n})}{\hat{N}(z,t_{n})+\epsilon}\right)+\epsilon\right),$
where, from Algorithm 1:
$\hat{R}(z,t_{n})=\sum_{i=1}^{n}r(x_{t_{i}})\exp{\left(-\sum_{j=1}^{d}\frac{(z_{j}-z_{j}(x_{t_{i}}))^{2}}{2{\sigma}_{j}^{2}}\right)},$ (8)
$\hat{N}(z,t_{n})=\sum_{i=1}^{n}\exp{\left(-\sum_{j=1}^{d}\frac{(z_{j}-z_{j}(x_{t_{i}}))^{2}}{2{\sigma}_{j}^{2}}\right)}.$ (9)
Since the domain is bounded, we know that for fixed $z$, both (8) and (9) have limit at infinity, i.e. $\lim_{n\rightarrow\infty}\hat{R}(z,t_{n})=\infty$ and $\lim_{n\rightarrow\infty}\hat{N}(z,t_{n})=\infty$. Hence, by Lemma C.1, we have
$\lim_{n\rightarrow\infty}\frac{\hat{R}(z,t_{n})}{\hat{N}(z,t_{n})+\epsilon}=\lim_{n\rightarrow\infty}\frac{\hat{R}(z,t_{n})}{\hat{N}(z,t_{n})},$
@@
$\lim_{n\rightarrow\infty}\frac{\hat{R}(z,t_{n})}{\hat{N}(z,t_{n})}=\lim_{n\rightarrow\infty}\frac{\sum_{i=1}^{n}r(x_{t_{i}})\exp\left(-\sum_{j=1}^{d}\frac{(z_{j}-z_{j}(x_{t_{i}}))^{2}}{2{\sigma}_{j}^{2}}\right)}{\sum_{i=1}^{n}\exp\left(-\sum_{j=1}^{d}\frac{(z_{j}-z_{j}(x_{t_{i}}))^{2}}{2{\sigma}_{j}^{2}}\right)}.$
Recall that metadynamics eventually leads to uniform sampling over the domain, independent of the potential. Hence, since $R$ and $z(x)$ are analytic, by Lemma C.2 we may replace the metadynamics samples $x_{t_{i}}$ with samples $u_{i}$ from a uniform distribution over $\mathcal{X}$:
$\lim_{n\rightarrow\infty}\frac{\hat{R}(z,t_{n})}{\hat{N}(z,t_{n})}=\lim_{n\rightarrow\infty}\frac{\sum_{i=1}^{n}r(u_{i})\exp\left(-\sum_{j=1}^{d}\frac{(z_{j}-z_{j}(u_{i}))^{2}}{2{\sigma}_{j}^{2}}\right)}{\sum_{i=1}^{n}\exp\left(-\sum_{j=1}^{d}\frac{(z_{j}-z_{j}(u_{i}))^{2}}{2{\sigma}_{j}^{2}}\right)}.$
@@
$\lim_{n\rightarrow\infty}\frac{\hat{R}(z,t_{n})}{\hat{N}(z,t_{n})}=\frac{\int_{\mathcal{X}}r(x^{\prime})\exp\left(-\sum_{j=1}^{d}\frac{(z_{j}-z_{j}(x^{\prime}))^{2}}{2{\sigma}_{j}^{2}}\right)\mathrm{d}x^{\prime}}{\int_{\mathcal{X}}\exp\left(-\sum_{j=1}^{d}\frac{(z_{j}-z_{j}(x^{\prime}))^{2}}{2{\sigma}_{j}^{2}}\right)\mathrm{d}x^{\prime}}.$
The limit is therefore a (scaled) convolution of the reward function with a Gaussian in the collective variable space with width vector $\sigma$. Taking the limit $\sigma_{j}\rightarrow 0$ for all $j\in\left\{1,2,\dots,d\right\}$, the Gaussian convergences to a delta distribution in the collective variable space and we have
$\lim_{\sigma\rightarrow 0}\lim_{n\rightarrow\infty}\frac{\hat{R}(z,t_{n})}{\hat{N}(z,t_{n})}=\int_{\mathcal{X}}r(x^{\prime})\delta(z-z(x^{\prime}))\mathrm{d}x^{\prime}.$
@@
$\lim_{\epsilon\rightarrow 0}\lim_{\sigma\rightarrow 0}\lim_{n\rightarrow\infty}\hat{V}(z,t_{n})$ $=-\frac{1}{\beta^{\prime}}\lim_{\epsilon\rightarrow 0}\log\left(\int_{\mathcal{X}}r(x^{\prime})\delta(z-z(x^{\prime}))\mathrm{d}x^{\prime}+\epsilon\right)$ (10)
$=-\frac{1}{\beta^{\prime}}\int_{\mathcal{X}}\log\left(r(x^{\prime})\right)\delta(z-z(x^{\prime}))\mathrm{d}x^{\prime}$ (11)
$=\int_{\mathcal{X}}V(x^{\prime})\delta(z-z(x^{\prime}))\mathrm{d}x^{\prime}\vcentcolon=V(z),$ (12)
@@
Appendix D Experiment details
D.1 Common experimental parameters
Training parameters: For all environments we train for $B=10^{5}$ batches use a learning rate with a linear schedule, starting at $10^{-3}$ and finishing at 00. For the TB loss, we train the $\log Z_{\theta}$ with a higher initial learning rate of $10^{-1}$ (also linearly scheduled). For the STB loss, we use $\lambda=0.9$, a value that has worked well in the discrete setting [15]. We use the Adam optimiser with gradient clipping. For all loss functions, we clip the minimum log-reward signal at $-10$. This enables the model to learn despite regions of near-zero reward between the modes of $r(x)$.
Replay buffer: We use a replay buffer with capacity for $10^{4}$ trajectories. Trajectories are stored in the replay buffer only if the terminal state’s reward exceeds $10^{-3}$. When drawing a replay buffer batch, trajectories are bias-sampled: $50\%$ randomly drawn from the upper $30\%$ of trajectories with the highest rewards, and the remaining $50\%$ randomly drawn from the lower $70\%$.
@@ D.1 Common experimental parameters
Noisy exploration: An additional constant, $\bar{\sigma}$, is added to the standard deviations of the Gaussian distributions of the forward and backward policies. Specifically, the forward policy becomes $\hat{p}_{F}(s_{t},s_{t-1};\theta)=\sum_{i=1}^{k}w_{i}\mathcal{N}(\mu_{i},(\sigma_{i}+\bar{\sigma})^{2})$, and similarly for the backward policy, where $k$ is the number of Gaussians in the mixture. We schedule the value of $\bar{\sigma}$ so that it decreases during training according to an exponential-flat schedule:
@@
where $j\in(1,\dots,B)$ is the batch number and $\bar{\sigma}_{0}=2$ is the initial noise, plotted in Figure 7. Note that, for the alanine dipeptide environment, the policy is a mixture of bivariate von Mises distributions and the noise $\bar{\sigma}$ is added to the concentration parameter $\kappa$, where concentration is related to standard deviation through $\sigma=\frac{1}{\kappa^{2}}$.
Thompson sampling: We use $10$ heads with the bootstrapping probability parameter set to $p=0.3$.
Local Search GFN: Local search samples trajectories using a forward-backward reconstruction method. First, forward policy samples actions to generate batch size $b$ on-policy trajectories. Then, the backward policy is used to re-sample the last $K$ steps of each trajectory. Finally, the forward policy is used to reconstruct the last $K$ steps, and trajectories in the batch with higher rewards from reconstruction replace the originals. We use $K=1$ for the line environment, alanine dipeptide environment and two-dimensional grid. We use $K=2$ and $K=3$ for the three and four dimensional (hyper)grids, respectively.
Nested sampling: We use the Gradient Based Nested Sampling (GBNS) [34] with a publicly available implementation: https://github.com/Pablo-Lemos/GGNS. For each environment, GBNS generates samples from the reward distribution. During training, we alternate between two types of backpropagation: one fully on-policy, and the other using backward-sampled trajectories from the generated samples, guided by the current backward policy.
Compute resources: Experiments are performed in PyTorch on a desktop with 32Gb of RAM and a 12-core 2.60GHz i7-10750H CPU.
D.2 Line environment details
Parameterisation: The forward and backward policies are a mixture of three Gaussians. We parameterise $\hat{p}_{F}$, $\hat{p}_{B}$ and the flow $\hat{f}$ through an MLP with 3 hidden layers, 256 hidden units per layer. We use the GELU activation function and dropout probability $0.2$ after each layer. This defines the torso of the MLP. Connecting from this common torso, the MLP has three single-layer, fully-connected heads. The first two heads have output dimension $9$ and parameterise the $3$ means ($\mu$), standard deviations ($\sigma$) and weights ($w$) of the mixture of Gaussians for the forward and backward policies respectively. The third head has output dimension $1$ and parameterises the flow function $\hat{f}$. The mean and standard deviation outputs are passed through a sigmoid function and transformed so that they map to the ranges $\mu\in(-14,14)$ and $\sigma\in(0.1,1)$. The mixture weights are normalised with the softmax function. The exception to this parameterisation is the backward transition to the source state which is fixed to be the Dirac delta distribution centred on the source, i.e. $\hat{p}_{B}(s_{0}|s_{1};\theta)=\delta_{s_{0}}$. For the TB loss, we treat $\log Z_{\theta}$ as a separate learnable parameter. We use batch size $b=64$. It takes approximately $1$ hour to train a continuous GFlowNet in this environment with $B=10^{5}$ batches.
@@ D.2 Line environment details
MetaGFN: We use $\Delta t=0.05$, $n=2$, $\beta=1$, $\gamma=2$, $w=0.15$, $\sigma=0.1$, $\epsilon=10^{-3}$. The domain of Adapted Metadynamics is restricted to $[-5,23]$ and reflection conditions are imposed at the boundary. The bias and KDE potentials are stored on a uniform grid with grid spacing $0.01$. Initial metadynamics samples are drawn from a Gaussian distribution, mean 00 and variance $1$, and initial momenta from a Gaussian distribution, mean 00 and variance $0.5$.
@@
Evaluation: The L1 error between the known reward distribution, $\rho(x)=r(x)/Z$, and the empirical on-policy distribution, denoted $\hat{\rho}(x)$, estimated by sampling $10^{4}$ on-policy trajectories and computing the empirical distribution over terminal states. Specifically, we compute $\text{error}=\frac{1}{2}\int_{-5}^{23}\left|\hat{\rho}(x)-\frac{r(x)}{Z}\right|\mathrm{d}x$, where the integral is estimated by a discrete sum with grid spacing $0.01$. Note that this error is normalised such that for all valid probability distributions $\hat{\rho}(x)$, we have $0\leq\text{error}\leq 1$.
On-policy distributions: Figure 9 shows the forward policy and replay buffer distributions (with bias sampling) after training for $10^{5}$ iterations with TB loss. MetaGFN is the only method that can uniformly populate the replay buffer and consistently learn all three peaks.
Adapted Metadynamics: Figure 10(a) shows the L1 error between the density implied by the KDE potential and the true reward distribution during a typical training run. Figure 10(b) shows the resulting $\hat{V}$ and $V_{\text{bias}}$ at the end of the training. By (1), Adapted Metadynamics has fully explored the central peaks. At (2), the third peak is discovered, prompting rapid adjustment of the KDE potential. By $2.5\times 10^{4}$ iterations, a steady state is reached and the algorithm is sampling the domain uniformly.
Sensitivity analysis of hyperparameters: We compare MetaGFN performance for various freqRB and freqMD values, showing the mean and standard deviation of the final L1 error after $10^{5}$ batches (Figure 8). Intuitively, we expect that if freqRB is set too large, then MetaGFN focuses predominantly on the on-policy dynamics and may fail to learn to sample distant modes. By contrast, setting $\texttt{freqRB}=1$ means exclusively training on the replay buffer, without any on-policy training trajectories, which we would expect to lead to slow convergence. Therefore, an intermediate value of freqRB should be optimal. We found that setting $\texttt{freqRB}=2$ (giving an equal weighting to both on-policy and replay buffer trajectories) showed the lowest loss. By contrast, performance is not too sensitive to freqMD. freqMD should be set so that Adapted Metadynamics explores the reward landscape in a comparable time it takes to train the GFN to convergence, hence only the order of magnitude of freqMD is important. For concreteness, we choose $\texttt{freqMD}=10$ but the model performance is not too sensitive to this parameter.
Figure 8: Mean and standard deviation L1 loss after $10^{5}$ training batches over $10$ repeats in the line environment. Blue line: changing freqRB while $\texttt{freqMD}=10$. Red line: changing freqMD while $\texttt{freqRB}=2$.
@@
Comparing MetaGFN variants: We consider three MetaGFN variants. The first variant, always backwards sample, regenerates the entire trajectory using the current backward policy when pulling from the replay buffer. The second variant, reuse initial backwards sample, generates the trajectory when first added to the replay buffer and reuses the entire trajectory if subsequently sampled. The third variant, with noise, is to always backwards sample with noisy exploration as per equation (13). We plot the L1 policy errors in Figure 11. We observe that always backward sample is better than reuse initial backwards sample for all loss functions. For DB and TB losses, there is no evidence for any benefit of adding noise, whereas noise improves training for STB loss, performing very similarly to TB loss without noise.
@@
D.3 Alanine dipeptide environment details
Figure 12: The potential KDE $\hat{V}$ learnt using MetaGFN on the alanine dipeptide environment, concentration $\kappa=10$.